OpenStudy (anonymous):

A 1.0 kg flywheel of radius 0.30m,initially at rest rotates 1000 times in 10.0 s. compute personally the following a)angular acceleration b)final angular velocity c)average linear velocity d) change in rotational kinetic energy

OpenStudy (michele_laino):

final angular speed is: \(\omega = 2\pi n\) where \(n=1000\)

OpenStudy (michele_laino):

so, angular acceleration is: \(\alpha = \frac{\omega }{{\Delta t}}\), where \(\Delta t=10.0\)

OpenStudy (michele_laino):

change in rotational kinetic energy is: \[\Delta KE = \frac{1}{2}I{\omega ^2}\] where \(I\) is the inertia momentum of the wheel

OpenStudy (michele_laino):

finally: average linear speed is: \[\frac{{0 + \omega r}}{2} = \frac{{\omega r}}{2}\] where \(r\) is the radius of the wheel

OpenStudy (anonymous):

Ina shoots a large marble (Marble A, mass: 0.08 kg) at a smaller marble (Marble B, mass: 0.05 kg) that is sitting still. Marble A was initially moving at a velocity of 0.5 m/s, but after the collision it has a velocity of -0.1 m/s. What is the resulting velocity of marble B after the collision? Be sure to show your work for solving this problem along with the final answer. (5 points)